OpenStudy (anonymous):

El Capitan in Yosemite National Park is 3593 ft high. How long would it take a carabiner to fall freely from the top?

OpenStudy (anonymous):

h=1/2at^2

OpenStudy (anonymous):

isnt there a formula you are supposed to use with the question ?

OpenStudy (anonymous):

need help with this

OpenStudy (anonymous):

t=sqrt(3593*2/32.2)=14.94 seconds